OpenStudy (anonymous):

Please Help I have a test tomorrow and I don't understand how to solve this :( Solve by factoring.. 4x^2 − 20 = 5

OpenStudy (anonymous):

i forgot about the 4, sorry

OpenStudy (anonymous):

\[4x^{2}-25=0\] use the quadratic equation to solve... x1=2.5 x2=-2.5

OpenStudy (anonymous):

i should shut up, two mistakes let me try one more time \[4x^2-25=0\] \[(2x+5)(2x-5)=0\]

OpenStudy (anonymous):

so \(2x+5=0\iff 2x=-5\iff x=-\frac{5}{2}\)

OpenStudy (anonymous):

and the other one is \[2x-5=0\iff 2x=5\iff x=\frac{5}{2}\]

OpenStudy (anonymous):

\[x=(-b \pm \sqrt{b ^{2}-4ac})/2a\] when a=4 b=0 c=-25

OpenStudy (anonymous):

you can also do this \[4x^2=25\] \[x^2=\frac{25}{4}\] \[x=\pm\sqrt{\frac{25}{4}}=\pm\frac{5}{2}\]

OpenStudy (anonymous):

Thank you guys